the mole fraction of  constituent A in vapour in equilibrium with binary liquid solution A and B is 3/4. if the ratio of vapour pressure of A and B at the same temperature is 3:1, what is the molefraction of A  in liquid phase?


a)1/2


b)1/4


c)1


d)3/4

Grade:12

1 Answers

Vikash Kumar
10 years ago

from Raoult''s law, the partial vapour pressure of each component in the solutionis directly proportional to its mole fraction.

so partial pressure of A= P0(A) * XA = (3/4)P (let) (as mole fraction is  directly prop. to partial pressure)

similarly, partial pressure of B= P0(B) * XB = (1/4)P 

putting P0(A) =3 and P0(B) = 1 we get XA = XB =1/2 (A)

Thanks & Regards
Vikash Kumar, 
askIITians Faculty
B.Tech, IIT Delhi